OpenStudy (anonymous):

simplafy (3/4)squared +1/9 divided by 1/3 showing all work

OpenStudy (anonymous):

\[(3/4)^2+\frac{1}{9}/\frac{1\text{ }}{3 } \longrightarrow \frac{9}{16}+\frac{1}{3} \longrightarrow \frac{43}{ 48} \]